OpenStudy (anonymous):

GIVING MEDAL; Find the angular displacement of a point P on a wheel for each of the given rotations. 1.400 2.7/12

OpenStudy (31356):

What do you think?

OpenStudy (anonymous):

I don't even know where to start, they're practice problems but she told us to give it a try :( @31356

OpenStudy (31356):

Oh okay, I am not sure either :( Good luck! Try asking other mods/ambassadors maybe for their help?

OpenStudy (anonymous):

Thanks, I will! @31356

OpenStudy (31356):

Good luck! :D

OpenStudy (ybarrap):

Can you do the next one?

OpenStudy (anonymous):

Kinda, but how would you start since it's a fraction? @ybarrap

OpenStudy (ybarrap):

If the rotation is 1.40 then it over rotated by 0.4 So $$ \cfrac{\theta}{2\pi}=0.4\\ \theta=0.4\times2\pi $$

OpenStudy (ybarrap):